OpenStudy (anonymous):

what is the ordered pair for y=-9x+6

OpenStudy (raden):

do u have choices

OpenStudy (anonymous):

(0,?)

OpenStudy (raden):

ohhh... just plug x=0, then solve for y

OpenStudy (raden):

y = -9x + 6 y = -9(0) + 6 y = ....

OpenStudy (anonymous):

y=6

OpenStudy (raden):

yes, correct so, the order pair is (0,6)

OpenStudy (raden):

actually it's called y-intercept